OpenStudy (anonymous):

Calculate the pH of 100 mL of a buffer solution containing 0.05 M benzoic acid (C6H5CO2H; Ka = 6.4 x 10-5) and 0.05 M sodium benzoate (NaC6H5CO2).(Note: please report your pH value to two decimal places.) ***I got this part*** (f) Calculate the pH after adding 10 mL of 0.10 M NaOH to the solution in part (d).

OpenStudy (aaronq):

so subtract the moles of acid neutralized by NaOH, and them to the moles of the conjugate base. Plug the values back into the Henderson-Hasselbalch equation.

OpenStudy (anonymous):

sorry but not sure what to do still

OpenStudy (aaronq):

how many moles of NaOH did you add? How many moles of NaC6H5CO2 and C6H5CO2H did you have before you added NaOH?

OpenStudy (anonymous):

.0001 moles NaOH? .005 moles of others?

OpenStudy (aaronq):

okay, so you neutralized some of the acid new moles of acid=0.005-0.0001 those moles became the conjugate base of the buffer new moles of base=0.005+0.0001 find new molarity (take the volume of NaOH into account) plug back into formula

OpenStudy (anonymous):

I got 12 and it says it's wrong. I am lost on this one.

OpenStudy (aaronq):

\(pH=pKa+log(\dfrac{(\dfrac{0.005+0.0001}{0.11L})}{(\dfrac{0.005-0.0001}{0.11L})}\)

OpenStudy (anonymous):

That gives me 4.28 which it counts wrong...

OpenStudy (anonymous):

rounded wrong it was 4.29. thanks!

OpenStudy (aaronq):

no problem !